OpenStudy (lukecrayonz):

Differentiate implicitly to find dy/dx, then find the slope at (2,5) x^2*y-4x^2-4=0

OpenStudy (lukecrayonz):

@Luigi0210 I got (-2y/x)+(8/x) but what do I do for the slope?:(

OpenStudy (loser66):

i am sorry, that's not right. redo, please

OpenStudy (loser66):

\[(x^2y - 4x^2-4)'\neq -\frac{2y}{x}+\frac{8}{x}\]

OpenStudy (anonymous):

Multiply the equation by x^(-2) which is same as dividing by x^2 Then isolate y Later differentiate the given equation you will get dy/dx in terms of x...then from given information substitute x=2 You will get the value of dy/dx at (2,5) which is same as the required slope.

OpenStudy (loser66):

try to help, but where is the Asker?

OpenStudy (lukecrayonz):

How is it not -2y/x+8/x? I mean simplified down it's (8-2y)/x

OpenStudy (loser66):

oh yes, you are right, I am so sorry, my bad. so, you have y ' = 8/x - 2y/x just replace(x, y) = ( 2,5) you get y ' = -1 and that is the slope

OpenStudy (loser66):

@Lukecrayonz I am sorry for my wrongness.

OpenStudy (lukecrayonz):

It's okay! Haha

OpenStudy (loser66):

I shouldn't help others right now. I am too busy with my stuff, hihihi... I thought I can relax a little bit by solving the easy stuff but I was wroooooong. XD